# Improving Dialog Systems with Natural Language Feedback and Preferences

### v1

@@ 6.4.1 Results on Standard Metrics
Valid Test Test Unseen
[…]
Table 12: Case study for combining system-level and instance-level feedback: performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence, Sat.: satisfaction. The system trained with human feedback-guided refinement (middle row) is better than no feedback-guided refinement (first row) or GPT-3.5 feedback-guided refinement (last row).
Standard metrics are shown in Table 9. Overall, we make the following observations: (i) When using the BB2 response generator, the query generator we trained improves the quality of the final generated response compared to the BB2 query generator. (ii) Training the response generator on satisfied data leads to further improvements when using our best query generator. (iii) However, when training on both satisfied data and refinement data, there was no improvement over the case when we train on the satisfied data only.
The main reason for observation (iii) is due to the gold response collection process of the FITS dataset. Most of the time, the gold response is a human verified response that was actually generated by the original BB2 with a few rounds of guidance from the user. Therefore, using standard reference-based metrics will favor responses that come from the original BB2 distribution. In addition, low-quality references will lead to underestimates of model performance when using reference-based metrics (Zhang et al., 2023a).
6.4.2 Results on Our Designed Metrics
The results when using our criteria-based metrics are shown in Table 10. Overall, we make the following observations: (i) When using BB2 response generator, our trained query generator helps improve the quality of the final generated responses from all perspectives compared to the BB2 query generator. (ii) When equipped with our trained query generator, training the response generator on satisfied data leads to consistent improvements of the helpfulness metric compared to the BB2 response generator, which indicates the importance of domain-adapted training. (iii) Training the response generator on both satisfied data and refinement data improves the quality of final responses from all perspectives compared to training on satisfied data only, demonstrating the importance of refinement data, which provides valuable guidance on the situations where models are prone to make mistakes. (iv) In terms of satisfaction, the best overall performing system is the one that consists of our query generator and our response generator, both trained on satisfied data and refinement data.
7 Combining System-level Feedback and Instance-level Feedback
Previous studies (Scheurer et al., 2022b; Shi et al., 2022; Chen et al., 2023a) have shown the effectiveness of instance-level feedback in the refinement process. This raises the question of whether we can take advantage of both system-level and instance-level feedback to improve the performance of dialogue systems further. We use response generation as a case study to explore the answer to this question. In particular, for the 1,000 unsatisfied responses we selected in §6.2.1, we collect their corresponding human feedback as well as GPT-3.5 generated feedback (prompts can be found in the Appendix) for the refinement process. We then design a text refinement prompt that combines both system-level feedback and instance-level feedback as shown in Table 11. We then introduce the following three systems for comparison.
[…]
Our system-level trained response generator using both satisfied and refinement data; equipped with our best query generator from §5. The system does not incorporate instance-level feedback.
[…]
Our system-level trained response generator equipped with trained query generator. Here, the response generator is trained on both satisfied data and our refinement data (where we incorporated human-written instance-level feedback into the response refinement prompt).
[…]
Our system-level trained response and query generators, where the response generator is trained on both satisfied data and our refinement data. Here, we incorporate GPT-3.5, rather than human, generated instance-level feedback into the response refinement prompt.
7.1 Results of Adding Instance-level Feedback
The results when using our designed metrics are shown in Table 12. We observe that adding human-written feedback to the response refinement part brings improvements in the five criteria-based metrics most of the time, and increases the overall satisfaction consistently. However, adding GPT-3.5 generated feedback results in consistent degraded performance in groundedness, factuality and confidence. Those observations raise two questions: (1) What are the differences between GPT-3.5 generated feedback and human-written feedback? (2) How does (human/GPT-3.5) feedback impact the refinement of responses? We answer those questions in turn in the following sections.
7.2 Comparison of Human and GPT-3.5 Feedback
Groundedness Factuality Helpfulness Relevance Confidence Satisfaction
No feedback refinement 39.16 90.35 83.48 98.10 100.00 76.50
H-FB refinement 40.11 87.50 81.10 97.80 99.84 74.60
GPT-3.5-FB refinement 32.77 81.50 90.20 98.40 99.84 79.50
Table 13: Combining system-level and instance-level feedback: the quality of refinements obtained using no/human/GPT-3.5 feedback. Human feedback-guided refinement (middle row) is more grounded and factual, which end up being important factors for final model training, see §7.3.
It is interesting to investigate the differences between human-written feedback and GPT-3.5 generated feedback to understand why adding human feedback is more helpful than GPT-3.5 feedback. In particular, we characterize feedback from the following quantifiable perspectives.
Refinement Success Rate
measures whether the feedback is constructive for refining the bad response by calculating the percentage of refinements that are classified as satisfactory.
Verbosity
measures how verbose the feedback is by counting the average number of words within the feedback text.
Diversity
measures how diverse the language use is within a set of feedback. We use the percentage of unique words within the whole set of feedback to measure diversity.
Grammar
measures whether the feedback is grammatically correct. We use the percentage of feedback sentences that are grammatically correct to measure language proficiency.999 We use Gramformer for grammar error checking: https://github.com/PrithivirajDamodaran/Gramformer.
Figure 2: Characteristics of human-written and GPT-3.5 generated feedback in terms of refinement success rate, verbosity, diversity and grammar.
The characteristics of human-written and GPT-3.5 generated feedback are shown in Figure 2. Overall, we found that GPT-3.5 generates feedback text that is often more verbose with fewer grammar errors. However, the low language diversity of GPT-3.5 feedback also demonstrates that the feedback is too generic and not as targeted as human feedback. Through manual inspection, we also found that GPT-3.5 feedback tends to point out problems w.r.t. each criterion in a general tone, while human feedback is typically straightforward and to the point. Examples of GPT-3.5 feedback and human feedback are shown in the Appendix.
7.3 Feedback Impact on Refinements
While GPT-3.5 feedback generally leads to more frequent satisfaction with refinements (as indicated in Figure 2), the performance of the resulting dialog system trained with these refinements falls short w.r.t. all our designed metrics compared to the system trained using human feedback-driven refinements as shown in Table 12. This disparity merits additional investigation, emphasizing the need to delve deeper into the refinement characteristics beyond surface-level satisfaction. Therefore, we also evaluate the refinement quality using the designed fine-grained metrics from §6.1.1. The results are shown in Table 13. Refinements obtained using human feedback mainly stand out in terms of groundedness and factuality, which is in line with the feedback clusters in Table 6 where over 40% of the feedback suggests the bot focus more on the search results; that is, focusing more on the search results will make the refinements more grounded, which will lead to a more grounded final system (see results in Table 12). Since language models are known to hallucinate regardless of their size (Ji et al., 2023; Li et al., 2023), grounding their generations to the documents is important to ensure factuality. Hence, groundedness of refinements can play an essential role in the performance of trained models.
7.4 Advantages of Human Feedback
Overall, we find that feedback written by humans captures the main issues better than feedback written by GPT-3.5. For example, when a response does not answer a question, GPT-3.5 will say that the response is unhelpful because it does not contain the information the user wants. In contrast, feedback given by humans tends to tell the model the information in the search results that the user needs or gives additional hints so that the model can be guided to better generate the response. Therefore, although GPT-3.5 can generate seemingly informative feedback, it cannot replace human annotators at the moment.
8 Conclusion
In this paper, we have introduced a general framework for system-level use of natural language feedback. This is achieved by using a large set of instance-level feedback to derive a small set of criteria, that can be used for refinement prompt engineering and metric design. We demonstrate the effectiveness of system-level feedback by conducting two case studies, in generating queries and formulating dialogue responses. We further combine the system-level feedback and instance-level feedback in the refinement data construction process, and observe that the resulting trained response generator makes considerable improvements compared to using system-level feedback only. Finally, we explore the possibility of substituting instance-level human feedback with GPT-3.5 generated feedback. We found that human feedback stands out in its ability to capture the main issues, while GPT-3.5 generated feedback is lengthy and less focused. Future work could explore how to use this analysis to improve model-based feedback further to try to reach human level.
Appendix A Appendix
A.1 Evaluation with GPT-3.5
Since previous studies have demonstrated GPT-3’s reliability in the evaluation of aspects such as factuality (Luo et al., 2023), helpfulness (Fu et al., 2023), relevance (Fu et al., 2023), etc. We use GPT-3.5 to evaluate the following perspectives with the help of ChatEval (Yuan et al., 2023).
Query Specificity
@@ Appendix A Appendix
In particular, we use the prompt as shown in Table 14. Before applying it to measure the quality of query refinements. We manually labeled 50 search queries from the FITS training split, and each one’s specificity label was decided by three annotators through majority vote. We calculate the agreement between GPT-3.5 and human annotators, and the result is 80%.
@@
The prompt we use is shown in Table 15. We conducted a meta-evaluation where we manually labeled 50 responses from the FITS training split, and three annotators decided on each one’s factuality label through majority vote. The agreement between GPT-3.5 and human annotators is 88%.
@@
A.2 Feedback Generation with GPT-3.5
@@
A.3 Examples of GPT-3.5 and Human Instance-level Feedback

### v2

@@ A.1 Manual Efforts Required to Derive System-level Criteria
We use GPT-3.5 to measure specificity (Fu et al., 2023) where we concatenate the dialog context and search query together and ask GPT-3.5 to judge whether the search query is specific using the chain-of-thought technique (Wei et al., 2022). In particular, we use the prompt as shown in Table 16. Before applying it to measure the quality of query refinements. We manually labeled 50 search queries from the FITS training split, and each one’s specificity label was decided by three annotators through majority vote. We calculate the agreement between GPT-3.5 and human annotators, and the result is 80%.
@@
We concatenate the search documents and response, and ask GPT-3.5 to judge if all information in the response is supported by the search documents. The prompt we use is shown in Table 17. We conducted a meta-evaluation where we asked three NLP PhD students at the same university as the first author to manually label 50 responses from the FITS training split. The annotation guideline we showed them is the same as the prompt designed for GPT-3.5 evaluation. Then, the three annotators decided on each one’s factuality label through a majority vote. The agreement between GPT-3.5 and human annotators is 88%.
@@
A.4 Human Evaluation on Response Outputs
In §6.4, our analysis revealed that training the response generator with both satisfied data and refinement data does not yield superior performance over using satisfied data alone, as evidenced by the F1 score and Perplexity (PPL) metrics. We hypothesized that this outcome might be attributed to a data bias in the FITS dataset, wherein the gold standard references are frequently produced by the BB2 model. Consequently, standard reference-based metrics, such as F1, tend to favor responses that closely resemble BB2 outputs. This bias potentially results in the underestimation of performance for models generating responses deviating from the BB2 distribution.
[…]
Figure 3: Win rates for system trained with both satisfied and refinement data and system trained with satisfied data only.
A.5 Instance-level Feedback vs. System-level Feedback
We argue that one of the drawbacks of instance-level approaches that utilize NL feedback is that they typically assume that every instance receives a feedback text, which is not practical in the real world where feedback tends to be sparse. Therefore, we also conducted a comparison experiment that assumes sparse instance-level feedback. Specifically, we collected the first 200 unsatisfied responses into a sparse refinement training set, refined via instance-level feedback only. We then train the response generator on this set alongside the satisfied data and compare its performance to our system-level trained model. Table 20 shows the performance of the two models as measured by our designed metrics. The system-level trained response generator outperforms the sparse instance-level trained response generator by a large margin on all metrics, demonstrating the importance of system-level feedback in a sparse instance-level feedback setting.
Valid Test Test Unseen
[…]
A.6 Feedback Generation with GPT-3.5
@@ A.4 Human Evaluation on Response Outputs
A.7 Examples of GPT-3.5 and Human Instance-level Feedback